# A comprehensive study of the relations between the properties of planetary systems and the chemical compositions of their host stars

### v1

@@ IV.3 Abundance patterns for exoplanetary systems
In Section II.3, we discussed that the abundances for Sc I rely on a few lines and their errors are typically higher than for Sc II. For Cr, we have the opposite situation and the uncertainties for Cr II are larger on average. In Section III.2.1, we showed that the abundances of Ti I present a significant correlation with which is also seen in other works in the literature. In Figure 5, we can see that the dispersions for Sc I, Ti I and Cr II are larger relative to Sc II, Ti II and Cr I. Finally, the discussion in Section IV.2 showed that results are not always consistent for the two ionization stages. Thus, in the following discussions, we will adopt the more reliable abundances obtained from Sc II, Ti II and Cr I for Sc, Ti and Cr, respectively.
The analysis of individual planets might be affected by the multiplanetary systems since the abundances of their host stars are considered more than once in the statistical tests. To avoid this, here we divided our sample of 510 stars according to the classifications of planetary systems described in Section II.1. For the following tests, we compared the abundance distributions of the subsamples using the Cucconi test and considering that differences were statistically significant when -value 0.001. Table 4 summarizes the results of these comparisons and shows only the elements for which we found statistically significant differences in each comparison.
@@ IV.3.1 Planet Size, Orbital Period and Multiplicity
The first comparison we performed was between systems with only small (N = 472) and only large (N = 27) planets and we found that, as for Fe, the abundances [X/H] for all elements are systematically higher for the latter. In particular, our result for [Mg/H] agrees with the one found by lt25. The differences between the average abundances are shown in blue in Figure 10 and the mean and median values for all elements are 0.143 0.024 dex and 0.135 0.011 dex, respectively. This result is a direct consequence of the well-known planet-metallicity correlation, which is valid only for the overall population of large planets (e.g., fischer05; buchhave12; petigura18; ghezzi21; wilson22). The Cucconi test shows that the [Fe/H] distributions are statistically different and the same holds for the elements Al, Si, Sc, Ti, Cr and Co (see Table 4).
@@
We then compared systems with only hot (N = 241) and only warm (N = 192) planets and observed a statistically significant difference for Al, Si, Ca, Fe, Co and Ni (see Table 4). These results are consistent with the higher abundances found for stars with hot planets for iron (e.g., mulders16; petigura18; wilson18), as well as for other elements (wilson22), except for Mg and Mn. In Figure 10, we show the differences between the average abundances in red and it is clear that systems with hot planets are systematically more enriched in metals, with mean and median differences of 0.058 0.017 dex and 0.057 0.005 dex, respectively.
@@
The next comparison was between systems with only small hot (N = 223) and only small warm (N = 179) planets. We find that the subsamples are statistically different for Si, Fe and Co (see Table 4). The mean and median differences are 0.054 0.015 dex and 0.060 0.011 dex, respectively, which are very similar to the values found in the previous test (see also Figure 10), as expected due to the fact that our sample is mainly composed of systems with only small planets. For the comparison between systems with only large hot (N = 16) and only large warm (N = 10) planets, we found no statistically significant differences.
@@
The small sample sizes for the large planets prevent further divisions into subsamples. For this reason, we now take a closer look only at the sample of systems with small planets, dividing it into systems that contain only super-Earths (SE; N = 286) and sub-Neptunes (SN; N = 192). The comparison between systems hosting only SE and only SN yields no statistical differences for any of the elements. A similar result was found when comparing SE H (N = 163) and SE W (N = 40), SN H (N = 42) and SN W (N = 130), SE H and SN H, SE W and SN W and SE+SN H (N = 18) and SE+SN warm (N = 9) (see Table 4).
@@
The absence of differences is intriguing since wilson22 found distinct occurrence rates for hot and warm super-Earths as well as for hot and warm sub-Neptunes for all 10 elements analyzed, including seven elements in common with our work (Mg, Al, Si, Ca, Mn, Fe and Ni). However, we observe that SE H systems have typically larger abundances relative to SE W systems (mean and median differences are 0.097 0.023 dex and 0.099 0.014 dex, respectively). The SN H systems are also enriched relative to SN W systems, although the mean and median differences are smaller (0.050 0.017 dex and 0.047 0.020 dex, respectively). These results are consistent with the findings of wilson22 that differences in the occurrence rates of hot and warm planets are more pronounced for super-Earths relative to sub-Neptunes. Therefore, the fact that we did not observe any statistically significant differences in the Cucconi tests could result from different samples, classification of the planets (wilson22 uses the limit 4.0 R⊕ instead of 4.4 R⊕ to divide SN and SS), consideration of planetary systems as opposed to individual planets and the stellar abundances.
@@
We also considered the possible influence of multiplicity by comparing the following pairs of systems: small single (N = 330) versus small multi (N = 142), SE single (N = 176) versus SE multi (N = 45) and SN single (N = 154) versus SN multi (N = 32). We did not find any statistical differences for any of the elements (see Table 4), and these results are consistent with previous findings for Fe from weiss18 and ghezzi21 and for Mg from lt25. Finally, we considered simultaneously the radii and orbital periods of the planets as well as the multiplicity in the system with the following comparisons: SE single H (N = 140) versus SE single W (N = 36), SE multi H (N = 23) versus SE multi W (N = 4), SN single H (N = 38) versus SN single W (N = 116) and SN multi H (N = 4) versus SN multi W (N = 14). As before, we did not find statistically significant differences for any of the elements (see Table 4). It is interesting to note that the difference observed by ghezzi21 between SE single H and SE single W for Fe has not been recovered here (-value = 0.036). Although we use the same metallicities as in ghezzi21 for the studied stars, we recall that the sample discussed in that study was larger and included metallicities obtained by petigura17 for those stars having low signal-to-noise HIRES spectra or large projected rotational velocities.
@@
We also checked whether considering [X/Fe] (corrected or not for the trends with age) instead of [X/H] would change the results. Now we do not find any statistical differences in the comparisons between systems with small and large planets (consistent with the results for Mg from lt25) as well as small hot and small warm planets (see Table 4), supporting the conclusion that the differences observed for [X/H] simply reflect the planet-metallicity correlation. For the comparison between systems with only hot and only warm planets, we do not see statistical differences for all elements either and this result supports the conclusion that the differences found when using [X/H] are caused by the underlying influence of [Fe/H] in the occurrence rates of hot exoplanets. For the remaining comparisons, we see no statistical differences. In particular, the absence of a statistically significant difference in the comparison of [Mg/Fe] distributions for small single and small multi systems is consistent with the result found by lt25.
@@
Since the individual abundances did not reveal any clear patterns with the architectures of planetary systems, we also investigated them together using four different quantities: median [X/H], median [X/Fe], median [X/Fe] corrected for age trends and [Ref/H] = A(Ref)⋆ - A(Ref)⊙, where Ref means refractory elements, A(Ref) = 10A(X)), A(X)⋆ = [X/H]⋆ + A(X)⊙ and A(X)⊙ are the solar abundances for a given element X from asplund09. The results are very similar to those presented in Table 4 and we only found statistically significant differences for the median [X/H] in the comparisons Small - Large, Hot - Warm and Small H - Small W as well as for [Ref/H] in the first of these cases.
@@
Specific groups of elements have different roles in the planetary formation processes (e.g., hinkel19 and torres-quijano25). Namely, the lithophiles are found in rock-forming minerals important to planetary interiors, while the siderophiles combine with iron and are present in the planetary cores. We investigated if the abundances of these two groups correlate with the planetary architectures. Following hinkel19 and torres-quijano25, we considered the groups of lithophiles (Na, Mg, Al, Si, Ca, Sc, Ti, V and Mn) and siderophiles (Cr, Co and Ni) and, as for the refractories, we tested both their median and summed abundances. In all four cases, we found a significant difference between systems with only small and only large planets. For the comparisons Hot - Warm and Small H - Small W, there is a statistical difference for the median abundance of the lithophiles, which is not corroborated by their summed abundances.
@@
For completeness, we also performed the comparisons listed in Table 4 for [/Fe] and the abundance ratios Fe/Si and Mg/Si. We did not find any statistically significant differences, however the abundances of the -elements presented some interesting features that deserve a more detailed discussion.
IV.3.2 Low and high- systems
@@
Interestingly, the minimum [Fe/H] value is similar for low- and high- systems with only small planets. However, it is higher (by 0.16 dex) for low- systems with only large planets relative to their high- counterparts. In order to test the significance of this last result, we performed a Monte Carlo resampling of each stellar metallicity within its respective uncertainty considering both Gaussian and uniform distributions. For 10,000 realizations, the median differences are 0.15 0.02 dex and 0.16 0.01 dex, respectively. Therefore, the difference is statistically significant and it seems that higher abundances of the -elements are able to partially compensate for the lower iron abundances, allowing the formation of giant planets in somewhat more metal-poor environments. Although these results are consistent with the findings of adibekyan12a; adibekyan12b, the higher minimum metallicity for the high-[/Fe] systems with only large planets relative their counterparts with only small planets suggest that, even if the abundance of -elements is higher, the Fe abundance still acts as a limiting factor in the formation of large exoplanets. We note, however, that the Cucconi tests show no significant differences between the metallicity distributions of low- and high- systems with only large planets.
adibekyan12b explored the metallicity interval -0.65 [Fe/H] -0.3 and found that stars hosting exclusively small planets were typically more enhanced in -elements. Our sample has 20 systems in this low metallicity regime and all of them have only small planets. Since 13 of them (65%) belong to the low- sequence, our results support the conclusion that an enhancement in the overall metallicity is not a requirement for the formation of small planets (buchhave12). Note, however, that adibekyan12b used just the Ti abundance in his analysis while we adopted the average of Mg, Si and Ti (see Section IV.1). If we consider only the Ti abundance, 11 systems (55%) belong to the low- sequence, which confirms our previous conclusion.
@@ IV.4 Chemical abundance versus condensation temperature trends
In the previous section, we analyzed the individual abundances of each element and concluded that, in general, they closely follow the well-known planet-metallicity correlation. However, it has previously been suggested in the literature that abundance patterns for groups of elements could be related to the formation of different classes of planets (e.g., hinkel19; torres-quijano25). In particular, the slopes of the abundances of refractory elements (T 900 K) as a function of condensation temperatures could result from the formation of terrestrial or giant planets (e.g., melendez09; bedell18; booth20). However, more recent studies discuss that this chemical signature that was originally found for the Sun does not seem to be related to planetary formation (liu20; rampalli24; carlos25; rampalli25; see also the review by gustafsson25).
@@ IV.4.1 Complete sample
Our sample has 510 stars hosting 726 planets, among which we have 472 systems with only small planets and 286 systems with only super-Earths. Therefore, this large and diverse sample offers a unique opportunity to test whether the formation of terrestrial planets imprints chemical signatures on solar-type stars. Using the derived abundances [X/Fe], the respective uncertainties determined in this work and the 50% condensation temperatures for the elements from lodders25, we performed weighted least squares fits (WLS) to the data and determined the slopes and associated statistics for each host star; these are presented in Table 1. Here, we use only [X/Fe] abundances (instead of [X/H]) to minimize the effects of atomic diffusion due to the different ages and masses of stars (e.g., dotter17; bedell18).
@@
We also analyzed if there were correlations between slopes and the different types of planetary systems using the Cucconi non-parametric test to evaluate possibly significant differences (-value 0.001) between the distributions. We computed median slopes for the same pairs of subsamples listed in Table 4. Although all of these tests revealed no significant differences between the subsamples considered, we notice that in general systems with only hot planets have smaller median slopes than systems with only warm planets. This can be seen in the right panel of Figure 12, which shows the distributions of slopes for each subsample and is sorted in order of ascending median slope (red horizontal lines). With the exception of SN multi W systems, the left part of the figure (lower median slopes) is populated with stars that host only hot exoplanets.
@@
As an additional test, we determined the average [X/Fe] abundances for our entire sample of 510 stars and performed a weighted least squares fit as a function of the condensation temperatures of the elements. We show the fit in the upper left panel of Figure 13 and the calculated slope is (6 3) dex K-1, with -value = 1.630 and -value = 0.131. Thus, it is consistent with zero and the Spearman and Pearson tests show that the correlation is not significant.
@@
As previously discussed in the literature (e.g., schuler15; bedell18; liu20; rampalli25), the trends between chemical abundances and condensation temperatures of elements might be affected by Galactic Chemical Evolution (GCE). In this context, we tested correcting our [X/Fe] abundances for the correlation with ages using linear relations and the coefficients in Table 3 (but see the caveats of this assumption as discussed by rampalli25). We assume that the corrected abundances are given by [X/Fe]corr = [X/Fe] - (aX Age) + bX, where X is a given element. Using [X/Fe]corr, we repeated the above tests and found some interesting differences. First, there is a displacement of the slope distribution towards more negative values. The median slope is (0 10) dex K-1 and 50% of our stars are more refractory depleted relative to the Sun (i.e., have negative slopes), thus placing it as a typical star (see also rampalli25). Moreover, the slope for the average abundances of the entire sample decreases to (0 1) dex K-1 and it is still statistically consistent with zero. Therefore, these results reinforce that the Sun has a similar or lower content of refractory elements relative to our sample of planet-hosting stars.
@@
melendez09 found a break in the slopes of the abundances versus condensation temperatures for the Sun compared to those of the solar twins at 1200 K and this limit was also used by liu20 to separate volatile and refractory elements. If, for example, we simply remove the elements Na, Mn and Cu, which have 1200 K, and recompute the slopes for our sample, for the [X/Fe] abundances, the median slope increases to (10 17) dex K-1 and there is a larger percentage of stars with positive slopes (74%). However, we still do not find significant differences between the slope distributions for the different classes of exoplanetary systems. For [X/Fe]corr, there are negligible changes in the median slope and the percentage of positive slopes.
@@
Since the refractory depletion in the Sun has been proposed in the literature to be caused by the formation of terrestrial planets (melendez09), we also investigated the slopes obtained when using the average [X/Fe] abundances for systems with at least one super-Earth (N = 291; left middle panel of Figure 13) or super-Earths only (N = 221; right middle panel of Figure 13). We determined the values (5 3) dex K-1 for host stars having at least one super-Earth and (6 3) dex K-1 for host stars having super-Earths only, both of which have corresponding -values . The Pearson and Spearman tests also show that the correlations are not significant. We note that the slopes are also consistent with zero when the abundances corrected for age trends are used.
@@
We also computed the slopes of systems with at least one (N = 42) or only (N = 35) Earth-sized planets () and found positive values for both: (12 3) dex K-1 and (10 4) dex K-1, respectively. However, the -values for the slopes and the Pearson and Spearman tests are equal to or larger than 0.001. Using [X/Fe]corr, we find slopes different than zero within uncertainties and the associated -value is lower than 0.001 only for the systems with at least one Earth-sized planet. However, the Pearson and Spearman tests return -values 0.001 for both cases. In summary, these comparisons suggest that the Sun has, at most, the same amount of refractories relative to stars hosting Earth-sized planets as well as other planets (which are all sub-Neptunes for this subsample) and stars hosting only Earth-sized planets. Since the Sun is typically more depleted in refractories even when compared with stars that host Earth-sized planets, this suggests its chemical peculiarity might be explained by factors other than planet formation.

### v2

@@ IV.3 Abundance patterns for exoplanetary systems
In Section II.3, we discussed that the abundances for Sc I rely on a few lines and their errors are typically higher than for Sc II. For Cr, we have the opposite situation, and the uncertainties for Cr II are larger on average. In Section III.2.1, we showed that the abundances of Ti I present a significant correlation with , which is also seen in other works in the literature. In Figure 5, we can see that the dispersions for Sc I, Ti I, and Cr II are larger relative to Sc II, Ti II, and Cr I. Finally, the discussion in Section IV.2 showed that results are not always consistent for the two ionization stages. Thus, in the following discussions, we will adopt the more reliable abundances obtained from Sc II, Ti II, and Cr I for Sc, Ti, and Cr, respectively.
The analysis of individual planets might be affected by the multiplanetary systems since the abundances of their host stars are considered more than once in the statistical tests. To avoid this, here, we divided our sample of 510 stars according to the classifications of planetary systems described in Section II.1. For the following tests, we compared the abundance distributions of the subsamples using the Cucconi test and considering that differences were statistically significant when -value 0.001. Table 4 summarizes the results of these comparisons and shows only the elements for which we found statistically significant differences in each comparison.
@@ IV.3.1 Planet Size, Orbital Period and Multiplicity
The first comparison we performed was between systems with only small (N = 472) and only large (N = 27) planets, and we found that, as for Fe, the abundances [X/H] for all elements are systematically higher for the latter. In particular, our result for [Mg/H] agrees with the one found by Loaiza-Tacuri et al. (2025). The differences between the average abundances are shown in blue in Figure 10, and the mean and median values for all elements are 0.143 0.024 dex and 0.135 0.011 dex, respectively. This result is a direct consequence of the well-known planet-metallicity correlation, which is valid only for the overall population of large planets (e.g., Fischer and Valenti, 2005; Buchhave et al., 2012; Petigura et al., 2018; Ghezzi et al., 2021; Wilson et al., 2022). The Cucconi test shows that the [Fe/H] distributions are statistically different and the same holds for the elements Al, Si, Sc, Ti, Cr, and Co (see Table 4).
@@
We then compared systems with only hot (N = 241) and only warm (N = 192) planets and observed a statistically significant difference for Al, Si, Ca, Fe, Co, and Ni (see Table 4). These results are consistent with the higher abundances found for stars with hot planets for iron (e.g., Mulders et al., 2016; Petigura et al., 2018; Wilson et al., 2018), as well as for other elements (Wilson et al., 2022), except for Mg and Mn. In Figure 10, we show the differences between the average abundances in red, and it is clear that systems with hot planets are systematically more enriched in metals, with mean and median differences of 0.058 0.017 dex and 0.057 0.005 dex, respectively.
@@
The next comparison was between systems with only small hot (N = 223) and only small warm (N = 179) planets. We find that the subsamples are statistically different for Si, Fe, and Co (see Table 4). The mean and median differences are 0.054 0.015 dex and 0.060 0.011 dex, respectively, which are very similar to the values found in the previous test (see also Figure 10), as expected due to the fact that our sample is mainly composed of systems with only small planets. For the comparison between systems with only large hot (N = 16) and only large warm (N = 10) planets, we found no statistically significant differences.
@@
The small sample sizes for the large planets prevent further divisions into subsamples. For this reason, we now take a closer look only at the sample of systems with small planets, dividing it into systems that contain only super-Earths (SE; N = 286) and sub-Neptunes (SN; N = 192). The comparison between systems hosting only SE and only SN yields no statistical differences for any of the elements. A similar result was found when comparing SE H (N = 163) and SE W (N = 40), SN H (N = 42) and SN W (N = 130), SE H and SN H, SE W and SN W, and SE+SN H (N = 18) and SE+SN warm (N = 9) (see Table 4).
@@
The absence of differences is intriguing since Wilson et al. (2022) found distinct occurrence rates for hot and warm super-Earths as well as for hot and warm sub-Neptunes for all 10 elements analyzed, including 7 elements in common with our work (Mg, Al, Si, Ca, Mn, Fe, and Ni). However, we observe that SE H systems have typically larger abundances relative to SE W systems (mean and median differences are 0.097 0.023 dex and 0.099 0.014 dex, respectively). The SN H systems are also enriched relative to SN W systems, although the mean and median differences are smaller (0.050 0.017 dex and 0.047 0.020 dex, respectively). These results are consistent with the findings of Wilson et al. (2022) that differences in the occurrence rates of hot and warm planets are more pronounced for super-Earths relative to sub-Neptunes. Therefore, the fact that we did not observe any statistically significant differences in the Cucconi tests could result from different samples, classification of the planets (Wilson et al. 2022 uses the limit 4.0 R⊕ instead of 4.4 R⊕ to divide SN and SS), consideration of planetary systems as opposed to individual planets, and the stellar abundances.
@@
We also considered the possible influence of multiplicity by comparing the following pairs of systems: small single (N = 330) versus small multi (N = 142), SE single (N = 176) versus SE multi (N = 45) and SN single (N = 154) versus SN multi (N = 32). We did not find any statistical differences for any of the elements (see Table 4), and these results are consistent with previous findings for Fe from Weiss et al. (2018) and Ghezzi et al. (2021) and for Mg from Loaiza-Tacuri et al. (2025). Finally, we considered simultaneously the radii and orbital periods of the planets as well as the multiplicity in the system with the following comparisons: SE single H (N = 140) versus SE single W (N = 36), SE multi H (N = 23) versus SE multi W (N = 4), SN single H (N = 38) versus SN single W (N = 116) and SN multi H (N = 4) versus SN multi W (N = 14). As before, we did not find statistically significant differences for any of the elements (see Table 4). It is interesting to note that the difference observed by Ghezzi et al. (2021) between SE single H and SE single W for Fe has not been recovered here (-value = 0.036). Although we use the same metallicities as in Ghezzi et al. (2021) for the studied stars, we recall that the sample discussed in that study was larger and included metallicities obtained by Petigura et al. (2017) for those stars having low signal-to-noise HIRES spectra or large projected rotational velocities.
@@
We also checked whether considering [X/Fe] (corrected or not for the trends with age) instead of [X/H] would change the results. Now, we do not find any statistical differences in the comparisons between systems with small and large planets (consistent with the results for Mg from Loaiza-Tacuri et al. 2025) as well as small hot and small warm planets (see Table 4), supporting the conclusion that the differences observed for [X/H] simply reflect the planet-metallicity correlation. For the comparison between systems with only hot and only warm planets, we do not see statistical differences for all elements either, and this result supports the conclusion that the differences found when using [X/H] are caused by the underlying influence of [Fe/H] in the occurrence rates of hot exoplanets. For the remaining comparisons, we see no statistical differences. In particular, the absence of a statistically significant difference in the comparison of [Mg/Fe] distributions for small single and small multi systems is consistent with the result found by Loaiza-Tacuri et al. (2025).
@@
Since the individual abundances did not reveal any clear patterns with the architectures of planetary systems, we also investigated them together using four different quantities: median [X/H], median [X/Fe], median [X/Fe] corrected for age trends and [Ref/H] = A(Ref)⋆ - A(Ref)⊙, where Ref means refractory elements, A(Ref) = 10A(X)), A(X)⋆ = [X/H]⋆ + A(X)⊙ and A(X)⊙ are the solar abundances for a given element X from Asplund et al. (2009). The results are very similar to those presented in Table 4, and we only found statistically significant differences for the median [X/H] in the comparisons Small - Large, Hot - Warm and Small H - Small W as well as for [Ref/H] in the first of these cases.
@@
Specific groups of elements have different roles in the planetary formation processes (e.g., Hinkel et al. 2019 and Torres-Quijano et al. 2025). Namely, the lithophiles are found in rock-forming minerals important to planetary interiors, while the siderophiles combine with iron and are present in the planetary cores. We investigated if the abundances of these two groups correlate with the planetary architectures. Following Hinkel et al. (2019) and Torres-Quijano et al. (2025), we considered the groups of lithophiles (Na, Mg, Al, Si, Ca, Sc, Ti, V, and Mn) and siderophiles (Cr, Co, and Ni), and, as for the refractories, we tested both their median and summed abundances. In all four cases, we found a significant difference between systems with only small and only large planets. For the comparisons Hot - Warm and Small H - Small W, there is a statistical difference for the median abundance of the lithophiles, which is not corroborated by their summed abundances.
@@
For completeness, we also performed the comparisons listed in Table 4 for [/Fe] and the abundance ratios Fe/Si and Mg/Si. We did not find any statistically significant differences; however, the abundances of the -elements presented some interesting features that deserve a more detailed discussion.
IV.3.2 Low- and high- systems
@@
Interestingly, the minimum [Fe/H] value is similar for low- and high- systems with only small planets. However, it is higher (by 0.16 dex) for low- systems with only large planets relative to their high- counterparts. In order to test the significance of this last result, we performed a Monte Carlo resampling of each stellar metallicity within its respective uncertainty considering both Gaussian and uniform distributions. For 10,000 realizations, the median differences are 0.15 0.02 dex and 0.16 0.01 dex, respectively. Therefore, the difference is statistically significant, and it seems that higher abundances of the -elements are able to partially compensate for the lower iron abundances, allowing the formation of giant planets in somewhat more metal-poor environments. Although these results are consistent with the findings of Adibekyan et al. (2012b, a), the higher minimum metallicity for the high-[/Fe] systems with only large planets relative to their counterparts with only small planets suggests that, even if the abundance of -elements is higher, the Fe abundance still acts as a limiting factor in the formation of large exoplanets. We note, however, that the Cucconi test shows no significant differences between the metallicity distributions of low- and high- systems with only large planets.
Adibekyan et al. (2012a) explored the metallicity interval -0.65 [Fe/H] -0.30 and found that stars hosting exclusively small planets were typically more enhanced in -elements. Our sample has 20 systems in this low-metallicity regime, and all of them have only small planets. Since 13 of them (65%) belong to the low- sequence, our results support the conclusion that an enhancement in the overall metallicity is not a requirement for the formation of small planets (Buchhave et al., 2012). Note, however, that Adibekyan et al. (2012a) used just the Ti abundance in his analysis while we adopted the average of Mg, Si, and Ti (see Section IV.1). If we consider only the Ti abundance, 11 systems (55%) belong to the low- sequence, which confirms our previous conclusion.
@@ IV.4 Chemical abundance versus condensation temperature trends
In the previous section, we analyzed the individual abundances of each element and concluded that, in general, they closely follow the well-known planet-metallicity correlation. However, it has previously been suggested in the literature that abundance patterns for groups of elements could be related to the formation of different classes of planets (e.g., Hinkel et al., 2019; Torres-Quijano et al., 2025). In particular, the slopes of the abundances of refractory elements (T 900 K) as a function of condensation temperatures could result from the formation of terrestrial or giant planets (e.g., Meléndez et al., 2009; Bedell et al., 2018; Booth and Owen, 2020). However, more recent studies discuss that this chemical signature that was originally found for the Sun does not seem to be related to planetary formation (Liu et al. 2020; Rampalli et al. 2024; Carlos et al. 2025; Rampalli et al. 2025; see also the review by Gustafsson 2025).
@@ IV.4.1 Complete sample
Our sample has 510 stars hosting 726 planets, among which we have 472 systems with only small planets and 286 systems with only super-Earths. Therefore, this large and diverse sample offers a unique opportunity to test whether the formation of terrestrial planets imprints chemical signatures on solar-type stars. Using the derived abundances [X/Fe], the respective uncertainties determined in this work and the 50% condensation temperatures for the elements from Lodders et al. (2025), we performed weighted least squares fits (WLS) to the data and determined the slopes and associated statistics for each host star; these are presented in Table 1. Here, we use only [X/Fe] abundances (instead of [X/H]) to minimize the effects of atomic diffusion due to the different ages and masses of stars (e.g., Dotter et al., 2017; Bedell et al., 2018).
@@
We also analyzed if there were correlations between slopes and the different types of planetary systems using the Cucconi nonparametric test to evaluate possibly significant differences (-value 0.001) between the distributions. We computed median slopes for the same pairs of subsamples listed in Table 4. Although all of these tests revealed no significant differences between the subsamples considered, we notice that in general systems with only hot planets have smaller median slopes than systems with only warm planets. This can be seen in the right panel of Figure 12, which shows the distributions of slopes for each subsample and is sorted in order of ascending median slope (red horizontal lines). With the exception of SN multi W systems, the left part of the figure (lower median slopes) is populated with stars that host only hot exoplanets.
@@
As an additional test, we determined the average [X/Fe] abundances for our entire sample of 510 stars and performed a WLS fit as a function of the condensation temperatures of the elements. We show the fit in the upper left panel of Figure 13 and the calculated slope is (6 3) dex K-1, with -value = 1.630 and -value = 0.131. Thus, it is consistent with zero, and the Spearman and Pearson tests show that the correlation is not significant.
@@
As previously discussed in the literature (e.g., Schuler et al., 2015; Bedell et al., 2018; Liu et al., 2020; Rampalli et al., 2025), the trends between chemical abundances and condensation temperatures of elements might be affected by Galactic Chemical Evolution (GCE). In this context, we tested correcting our [X/Fe] abundances for the correlations with ages using linear relations and the coefficients in Table 3 (but see the caveats of this assumption as discussed by Rampalli et al. 2025). We assume that the corrected abundances are given by [X/Fe]corr = [X/Fe] - (aX Age) + bX, where X is a given element. Using [X/Fe]corr, we repeated the above tests and found some interesting differences. First, there is a displacement of the slope distribution toward more negative values. The median slope is (0 10) dex K-1, and 50% of our stars are more refractory depleted relative to the Sun (i.e., have negative slopes), thus placing it as a typical star (see also Rampalli et al. 2025). Moreover, the slope for the average abundances of the entire sample decreases to (0 1) dex K-1 and it is still statistically consistent with zero. Therefore, these results reinforce that the Sun has a similar or lower content of refractory elements relative to our sample of planet-hosting stars.
@@
Meléndez et al. (2009) found a break in the slopes of the abundances versus condensation temperatures for the Sun compared to those of the solar twins at 1200 K, and this limit was also used by Liu et al. (2020) to separate volatile and refractory elements. If, for example, we simply remove the elements Na, Mn and Cu, which have 1200 K, and recompute the slopes for our sample, for the [X/Fe] abundances, the median slope increases to (10 17) dex K-1, and there is a larger percentage of stars with positive slopes (74%). However, we still do not find significant differences between the slope distributions for the different classes of exoplanetary systems. For [X/Fe]corr, there are negligible changes in the median slope and the percentage of positive slopes.
@@
Since the refractory depletion in the Sun has been proposed in the literature to be caused by the formation of terrestrial planets (Meléndez et al., 2009), we also investigated the slopes obtained when using the average [X/Fe] abundances for systems with at least one super-Earth (N = 291; left middle panel of Figure 13) or super-Earths only (N = 221; right middle panel of Figure 13). We determined the values (5 3) dex K-1 for host stars having at least one super-Earth and (6 3) dex K-1 for host stars having super-Earths only, both of which have corresponding -values . The Pearson and Spearman tests also show that the correlations are not significant. We note that the slopes are also consistent with zero when the abundances corrected for age trends are used.
@@
We also computed the slopes of systems with at least one (N = 42) or only (N = 35) Earth-sized planets () and found positive values for both: (12 3) dex K-1 and (10 4) dex K-1, respectively. However, the -values for the slopes and the Pearson and Spearman tests are equal to or larger than 0.001. Using [X/Fe]corr, we find slopes different from zero within uncertainties, and the associated -value is lower than 0.001 only for the systems with at least one Earth-sized planet. However, the Pearson and Spearman tests return -values 0.001 for both cases. In summary, these comparisons suggest that the Sun has, at most, the same amount of refractories relative to stars hosting Earth-sized planets as well as other planets (which are all sub-Neptunes for this subsample) and stars hosting only Earth-sized planets. Since the Sun is typically more depleted in refractories even when compared with stars that host Earth-sized planets, this suggests its chemical peculiarity might be explained by factors other than planet formation.